OpenStudy (anonymous):

Which statement about subatomic particles is NOT true? A. Protons and neutrons have almost the same mass. B. Protons and electrons have opposite charges. C. Unlike protons and electrons, neutrons have no charge. D. Protons and neutrons have the same charge.

OpenStudy (maheshmeghwal9):

D is not true :D

OpenStudy (maheshmeghwal9):

Neutrons have no charge at all. Protons have \[+1.6 \times 10^{-19}C\]charge Electrons have \[-1.6 \times 10^{-19}C\]charge

OpenStudy (maheshmeghwal9):

Protons have mass equal to \[1.67 \times 10^{-27}kg.\]Electrons have mass equal to\[9.1 \times 10^{-31}kg.\] \[\color{blue}{\text{That's All.}}\]

OpenStudy (maheshmeghwal9):

one more thing Neutrons have mass equal to \[1.6 \times 10^{-27}kg.\]Same as proton.

OpenStudy (maheshmeghwal9):

These must be remembered these are very useful:D

OpenStudy (anonymous):

I NEEED A SIMPLE ANSWERRR :'(